OpenStudy (anonymous):

a net horizontal force of 4000N is applied to a 1400kg car. what will be the car's speed after 10seconds if it started from rest?

OpenStudy (espex):

Assuming constant acceleration: \[F=ma, v_f=v_i+at\]\[F=ma \rightarrow a=\frac{F}{m} \rightarrow a=\frac{4000N}{1400kg}=2.857\frac{m}{s^2}\] \[v_f=v_i+at \rightarrow v_f=0+(2.857\frac{m}{s^2} \times 10\sec)=28.57\frac{m}{s}\]